# Open-World Task and Motion Planning via Vision-Language Model Inferred Constraints

### v2

@@ V Experiments and System Demonstration
Figure 4 shows our main soundness rate results. These results demonstrate that OWL-TAMP generally does not produce false positives; in fact, there was only a single false positive across all tasks. This indicates that our approach is generally able to accurately capture the constraints within the various natural language goal expressions. By contrast several baselines — most notably ‘No VLM’ and ‘No cont.’ — have much higher satisfying plan rates than success rates (i.e., they falsely declare they have solved tasks). This is because they translate the natural language goal into problems that they solve, but that do not actually correspond to achieving the true underlying task goal.
Our approach failed at least once in ‘BerryCook’, ‘Mug3’, and ‘SoupPour’. Upon inspection, we found that all the failures in ‘BerryCook’ were due to incorrect continuous constraints. The VLM incorrectly constrained the strawberry to be simultaneously inside the bowl and pan at the end of the task in order to achieve the goal. In the ‘Mug3’ task, we found that our approach generated correct constraints, but was unable to find a satisfying set of samples with the provided budget. In the ‘SoupPour’ task, we found that the VLM generated incorrect discrete constraints: the TAMP system was not able to find a plan that included the generated partial plan as a sub-sequence.
Real-World System Deployment To test our approach on real-world hardware, We deployed OWL-TAMP on a custom dual-arm manipulator with two Kinova Gen3 arms and a pan-tilt head camera to perform a variety of natural language manipulation tasks. We demonstrated $19$ different tasks covering a wide range of generalized pick-and-place problems (i.e., each task was solvable via a specific pick-and-place sequence with appropriate continuous parameters) within this domain, $2$ of which are depicted in Figure 3. The goal descriptions for these tasks are listed below.
@@
Tasks were implemented by placing a set of objects in front of the robot and providing a natural language command. The robot leveraged a vision pipeline described in Appendix A-D to estimate the geometric scene, performed constraint generation and planning using this simulation, and then executed generated plans on the real hardware open-loop. We used the PDDLStream [9] system as the underlying TAMP system within OWL-TAMP. Execution videos of these tasks are included in our supplementary materials.
Overall, OWL-TAMP was successfully able to generate the necessary constraints to enable generally correct behavior in all these tasks, given just direct camera input and natural language goal descriptions. Importantly, we used the same system with no variation — even in prompting or the few-shot examples provided  [^4] We provided the same few-shot examples as used in our simulated tasks. — to accomplish this variety of tasks. Nearly all of the tasks required specific choices for continuous parameters (e.g. side grasps instead of purely top-down, constrained placements due to collisions and reachability), and several required the TAMP system to augment the VLM-proposed partial plan to move one or more objects out of the way. Several tasks included specific instructions that the TAMP system could not possibly have achieved given its internal predicates and operators and thus relied on the VLM-inferred constraints. Thus, our system’s ability to solve these tasks required close interplay between both the TAMP system and the VLM central to OWL-TAMP.
VI Limitations and Future Work
Our system relies entirely on the VLM to generate constraints that are both syntactically and semantically correct. There is currently no provision for recovering from errors — either syntactic or semantic — in the generated constraints. It is possible for the VLM to generate constraints — especially continuous constraints — for a task that are either not satisfiable at all, or that are very challenging to satisfy, which would cause the TAMP system to fail to find a solution. Finally, we assume a set of primitive parameterized robot skills (e.g. moving, rigid grasping, placement), which plans ultimately consist of. We also assume a library of primitive helper functions that the VLM leverages to construct continuous constraints, and we observed that generated constraints are usually simple compositions of helper functions.
Avenues for future work include enabling a feedback loop between the TAMP system and the VLM so that the VLM can modify or augment constraints if the TAMP system declares a problem infeasible, enabling the VLM to generate continuous parameter samplers to help the TAMP system satisfy generated continuous constraints, finetuning VLMs [58] on a representative distribution of robot manipulation tasks and constraints to improve their accuracy and generality at constraint generation, and investigating extensions of our framework to handling tasks with more complex (e.g. temporal) constraints.
VII Conclusion
We presented OWL-TAMP, a system that uses VLMs to infer discrete and continuous constraints that enable a TAMP system to interpret and accomplish open-world language instructions. We demonstrated that our approach supports multiple underlying TAMP systems, and is able to solve a wide range of complex, long-horizon manipulations tasks across different robot embodiments in simulation and the real world.
Acknowledgements
Appendix A Appendix
A-A Additional RAVENS-YCB Manipulation Task Details
[…]
  • $\bullet$
    Berry1: Goal description: “put the strawberry onto the light-grey region at the center of the table”. Objects: strawberry, light_grey_region, table_surface. The position of the light_grey_region is held constant at the center of the table. The optimal solution to this task requires a sequence of 2 skills.
  • $\bullet$
    Citrus: Goal description: “pack the citrus fruit onto the plate”. Objects: strawberry, apple, pear, lemon, orange, plum, table_surface. The plate object is always set to be at the center of the table in the initial state. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    Berry2: Goal description: “put the strawberry onto the light-grey region at the center of the table”. Objects: strawberry, light_grey_region, potted_meat_can, table_surface. The only object whose position is randomized in the initial state is the strawberry: the potted_meat_can is always set to totally obstruct the surface of the light_grey_region in the center of the table: it must be moved out of the way to make room to place the strawberry and successfully accomplish the goal description. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    BerryCook: Goal description: “Cook the strawberry by putting it in the pan, then finally simply place it in the bowl. The strawberry should only be in the bowl at the end!”. Objects: strawberry, skillet, bowl, table_surface. The only object whose position is randomized in the initial state is the strawberry. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    FruitSort: Goal description: “Put all the fruit to the left of the line bisecting the table”. Objects: pear, sugar_box, hammer, tomato_soup_can, strawberry, apple, red_line, table_surface. The position of the red_line object is kept constant at the center of the table. The optimal solution to this task requires a sequence of 6 skills.
  • $\bullet$
    Coffee: Goal description: “I want to pour some coffee into the cup; can you set up the cup on the table so I can do this properly?” (requires reorienting the cup so that it is placed ‘right-side-up’ on the table). Objects: mug, table_surface. The optimal solution to this task requires a sequence of 2 skills.
  • $\bullet$
    Mug1: Goal description: “Setup the mug so it’s upright, then put whatever object that fits inside of it”. Objects: fork, power_drill, potted_meat_can, mug, table_surface. The initial position of the mug is fixed. The optimal solution to this task requires a sequence of 4 skills.
  • $\bullet$
    Mug2: Goal description: “Place cutlery inside the mug and then place the mug itself on the table near the condiment” (the mug’s opening is obstructed by a large orange, which must be moved out of the way). Objects: fork, knife, sponge, strawberry, orange, mustard_bottle, mug, table_surface. The initial position of the mug and orange are fixed such that the orange is always obstructing the mouth of the mug. The optimal solution to this task requires a sequence of 8 skills.
  • $\bullet$
    Mug3: Goal description: “place cutlery into the mug and have the mug be near the condiment”. This is the same as CutleryInMug2, except in the initial state the mug is not obstructed by an orange, but rather has a ball stuck inside it. This must be removed by ‘pouring’ it out of the cup itself). Objects: fork, golf_ball, mustard_bottle, mug, table_surface. The initial position of the mug and golf_ball are fixed such that the white is always inside the mug. The optimal solution to this task requires a sequence of 8 skills.
  • $\bullet$
    SoupPour: Goal description: “Serve the fruits on the white mat (make sure the peach is to the right of the apple” and pour soup into the red container”. The white mat is originally obstructed by the soup can as well as a spam can, which must be moved out of the way to successfully place the fruits. Objects: white_mat, tomato_soup_can, potted_meat_can, bowl, apple, peach, table_surface. The only object positions that are randomized in the initial state are that of the apple and the peach. The optimal solution to this task requires a sequence of 10 skills.
A-B Additional Method Implementation Details
Here, we first provide a detailed explanation of how we implement the TAMP system used in our simulation experiments (Appendix  A). Subsequent sections (Appendix A-E, Appendix A-F, and Appendix A-G) provide more specific details around prompting. For details beyond those presented here, we refer the reader to our open-source code release, which we will make available after acceptance.
A-B1 TAMP System Details
We implement a relatively simple variant of the search-then-sample TAMP approach used in several recent works [8, 12, 16, 17] and use this as the underlying TAMP system for all methods that require one (i.e., all methods except ‘CaP’ and ‘CaP-sample’). We implement versions of the same predicates and operators described in Section III (including a version of transport that we call Pour), though we omit the move operator (movement is performed automatically as part of attach and detach) choose to name them differently (as illustrated by our prompts in Appendix A-F below). We manually define the initial state of all tasks in terms of literals involving these predicates; the robot always begins each task at the same initial configuration, and with the HandEmpty() predicate set to True. Additionally, we associate each operator with a natural language description to make it easier for the VLM to perform discrete constraint generation (e.g., for the detach operator, the description is something like: “places object o stably atop a surface; you can specify a description of how this placement should be performed by filling in the description parameter accordingly”). Each operator is linked to one particular low-level skill included with the environment.
We define an associated sampler for each of these operators. Each sampler takes in the current state of the task, as well as the operator’s discrete arguments, and defines a distribution over the operator’s continuous parameters. These samplers are setup to define broad distributions (e.g. the sampler for the detach operator simply tries to find a pose somewhere broadly above the surface to be detached onto, but roughly within the confines of the x and y boundaries of the surface’s axis-aligned bounding box), though we do modify these slightly depending on the task (e.g. the sampler for the ‘Coffee’ and ‘Mug’ tasks defines particular distributions on the orientations the mug and/or cutlery should be Placed).
Given these components, we can adopt the following hierarchical planning strategy to achieve a symbolic predicate goal $G$ from initial state $s_{0}$ 555see [12] or [16] for a fuller presentation of the search-then-sample TAMP strategy we employ.:
  1. 1.
  2. 2.
    For $b$ within the backtracking budget:
    1. (a)
      For each ground operator in this plan:
      1. i.
        If the preconditions do not hold, break.
      2. ii.
        For $i$ within the sampling budget:
        1. A.
          Call the associated sampler to get continuous parameters.
        2. B.
          Use these to instantiate the skill associated with the operator.
        3. C.
          Execute the skill within the planning model (i.e., the simulator)
        4. D.
          Check that the resulting state satisfies all the predicate effects.
Our task planning is performed initially via a simple $A^{*}$ search. During backtracking (i.e., when the sampling budget is exhausted for the first time and a new task plan is required), we employ a set of manually-engineered strategies to modify the task plan based on the most-recent failed operator (e.g. if the most recent-failed operator is a detach that was attempting to place an object atop a particular surface, and there are other objects atop that surface already, we randomly append a attach detach sequence to move one of those objects to a different part of the table).
A-B2 OWL-TAMP Constraint Generation Example Walkthrough
Below walk through a concrete example of generating both discrete and continuous constraints for the ‘Mug1’ task.
First, we prompt the model to generate discrete constraints given the task goal. It generates the following partial plan, along with language descriptions filled in for each action:
1(’Pick(mug)’, ’grasp the mug securely to lift it from the table.’)
3(’Pick(fork)’, ’grasp the fork securely to lift it from the table.’)
Next, we prompt the model to provide continuous constraints for the achieve_goal operator at the end of the plan. This corresponds to providing continuous constraints that correspond to the task’s goal description.
1Goal Check Functions:
2def goal_check0() -> bool:
4 return upright_mug
5
6
7def goal_check1() -> bool:
9 return position_within_bounds(fork.pose, inside_mug_bounds)
Next, we prompt the model to generate continuous constraints for each operator given these generated constraint functions as input. We start from the beginning of the plan. The Pick(mug) operator (which corresponds to attach) does not have any effects or constraint predicates that depend on the VLM, so we skip this operator. The Place_Ontop(mug, table) operator ((which corresponds to detach) does feature a VLMPose constraint. So we prompt the VLM to generate continuous constraints that implement it based on the language description as well as the goal generated constraints.
1Goal Check Functions:
2def goal_check2() -> bool:
The Place_Inside(fork, mug) operator (which also corresponds to detach) does not have any effects or constraint predicates that depend on the VLM, so we skip this operator as well. Finally, the Place_Inside(fork, mug) does have a VLMPose constraint, so we once again prompt the VLM, and obtain the following continuous constraints for this operator:
1Goal Check Functions:
2def goal_check2() -> bool:
4 return position_within_bounds(fork.pose, inside_mug_bounds)
Note that in this particular case, the VLM only checks that the fork is inside the mug, which is a constraint that is built into the TAMP system itself (via the Inside predicate that is an effect of the Place_Inside operator).
Now, we have a partial plan along with continuous constraints for particular actions. We call our underlying TAMP system to attempt to satisfy these, as well as its own internal constraints.
A-B3 Code as Policies Implementation Details
Following recent work [7], we adapt Code as Policies to our RAVENS YCB-Manipulation domain by having it output a Language Model Program (LMP) function that — given an initial state — directly outputs a full plan with skills and continuous parameters. We provide it with all the same helper functions used for continuous parameter generation that our method has access to. We also provide it with few-shot example solutions to the same 3 example problems as we provide in the continuous constraint generation prompt of our approach. Additionally, we provide it with one sampler per skill that it can use to generate continuous parameters for each respective skill. We provide the full detailed prompt we use for this method below in Appendix A-G.
Tasks
TABLE II: Number of samples on all tasks. We report the average number of continuous parameter sampling attempts for each task across $10$ random seeds; the number of samples within the 95% confidence interval is reported after the $\pm$ sign for every entry. We also include a row (‘Manual’) reporting results on a variant of OWL-TAMP where we manually specify ground-truth discrete and continuous constraints to serve as an approximate lower bound on the number of samples our approach could achieve.

### v4

@@ 6.2 Real-World System Deployment
7 Conclusion
We presented OWL-TAMP, a system that uses VLMs to generate discrete and continuous constraints that enable a base TAMP system to interpret and accomplish open-world language instructions. We demonstrated that OWL-TAMP solves a wide range of complex, long-horizon language-based manipulation tasks more reliably than existing baselines, and that it applies across different robot embodiments in simulation and the real world.
8 Limitations
Our system relies entirely on the VLM to generate constraints that are both syntactically and semantically correct. While the TAMP system is capable of backtracking internally, our current implementation cannot recover from errors in the generated constraints themselves. Another limitation is our assumption of a set of primitive parameterized robot controllers (e.g., moving, rigid grasping, placement) and a library of primitive helper functions — common in planning and foundation model-based code-generation works — which OWL-TAMP uses to generate plans and continuous constraints. Despite these limitations, our results highlight the promise of combining large-scale pretrained models with structured planners, and we hope this work inspires further progress toward flexible, capable robot manipulation systems.
Acknowledgements
Appendix A Appendix
We first provide details of how we implement OWL-TAMP (Appendix A.1), including details about its integration with standard TAMP systems, generation of discrete and continuous constraints. Subsequent sections provide more specific details around prompting the VLM (Appendix A.6A.7, and A.8). For details beyond those presented here, we refer the reader to our open-source code release, which we will make available after acceptance.
A.1 Additional Method Implementation Details
We provide a detailed explanation of how we implement the search-then-sample TAMP system used in our simulation experiments (Section 6.1).
A.1.1 Search-then-Sample TAMP System Details
In our simulation experiments, we implement a relatively simple variant of the search-then-sample TAMP approach used in several recent works [46, 35, 43, 34]. We use this as the underlying TAMP systems for all methods that require one (i.e., all methods except ‘CaP’ and ‘CaP-sample’). We implement versions of the same predicates and operators described in Section 3 (including a version of transport that we call Pour), though we omit the move operator as movement is performed automatically as part of attach and detach (as illustrated by our prompts in Appendix A.7 below).
We manually define the initial state of all tasks in terms of literals involving these predicates; the robot always begins each task at the same initial configuration, and with the HandEmpty() predicate set to True. Additionally, we associate each parameterized action with a natural language description to make it easier for the VLM to perform discrete constraint generation (e.g., for the detach operator, the description is something like: “places object o stably atop a surface; you can specify a description of how this placement should be performed by filling in the description parameter accordingly”). Each operator is linked to one particular low-level controller included with the RAVENS-YCB environment [53].
We define an associated sampler for each of these parameterized actions. Each sampler takes in the current state of the task, as well as the action’s discrete arguments, and defines a distribution over the action’s continuous parameters. These samplers are setup to define broad distributions (e.g.. the sampler for the detach operator simply tries to find a pose somewhere broadly above the surface to be detached onto, but roughly within the confines of the x and y boundaries of the surface’s axis-aligned bounding box), though we do modify these slightly depending on the task (e.g. the sampler for the ‘Coffee’ and ‘Mug’ tasks defines particular distributions on the orientations the mug and/or cutlery should be detach-ed).
Given these components, we can adopt the following hierarchical search-then-sample planning strategy to achieve a symbolic predicate goal $G$ (derived from a plan sketch via the procedure listed in Appendix A.1.2 from initial state $s_{0}$111see [35] or [43] for a fuller presentation of the search-then-sample TAMP strategy we employ.:
  1. 1.
  2. 2.
    For $b$ within the backtracking budget:
    1. (a)
      For each ground operator in this plan:
      1. i.
        If the preconditions do not hold, break.
      2. ii.
        For $i$ within the sampling budget:
        1. A.
          Call the associated sampler to get continuous parameters.
        2. B.
          Use these to instantiate the controller associated with the operator.
        3. C.
          Execute the controller within the planning model (i.e., the simulator)
        4. D.
          Check that the resulting state satisfies all the constraints, including manually-engineered and VLM-generated constraints.
We perform initial task planning via a simple $A^{*}$ search. During backtracking (i.e., when the sampling budget is exhausted for the first time and a new task plan is required), we employ a set of manually-engineered strategies to modify the task plan based on the most-recent failed operator (e.g. if the most recent-failed operator is a detach that was attempting to place an object atop a particular surface, and there are other objects atop that surface already, we randomly append a attach detach sequence to move one of those objects to a different part of the table).
A.1.2 Discrete Constraint Generation Details
Given a goal described in natural language $L$ (Section 3), we first prompt a VLM to generate a plan sketch (i.e., a partial plan) that serves as a discrete constraint on the space of TAMP solutions (Section 5.1). To enable this, we associate a natural language description of each available action with that particular action. Although we could directly prompt a VLM for relevant actions and goals, without a list of candidates, the VLM is likely to be syntactically and semantically inaccurate. Instead, we propose to first ground the set of reachable actions $A$ and literals $\Lambda$ (i.e., grounded predicates) available to the TAMP system before prompting the VLM to return values in these sets. We use relaxed planning [6, 7] from the initial state $s_{0}$ to simultaneously ground [24, 13, 23] and explore the sets of reachable actions $A$ and literals $\Lambda$. When instantiating continuous parameters, we use placeholder values, such as optimistic values [18, 17, 14, 19], to ensure a finite set of actions are instantiated. Similarly, we use placeholders for description parameters.
Algorithm 2 VLM Task Reasoning
1:procedure vlm-task-reasoning($s_{0},{\cal A},{\cal I},L$)
2:  $A\leftarrow\textsc{ground-actions}(s_{0},{\cal A})$
3:  $\Lambda\leftarrow s_{0}\cup\{l\mid a\in A.\;l\in e.{\it eff}\}$
[…]
5:  for $i\in[1,n-1]$ do
6:   $a_{i}.{\it eff}\leftarrow a_{i}.{\it eff}\cup\{{\texttt{Executed}}(i)\}$
7:   $a_{i+1}.{\it pre}\leftarrow a_{i+1}.{\it pre}\cup\{{\texttt{Executed}}(i)\}$
8:  $a_{n}.{\it eff}\leftarrow a_{n}.{\it eff}\cup\{{\texttt{Executed}}(n)\}$
9:  $G\leftarrow\{l_{1},..,l_{m}\}$
10:  return $\textsc{solve-tamp}(s_{0},A,G\cup\{{\texttt{Executed}}(n)\})$
Algorithm 2 presents the VLM partial plan generation pseudocode. It takes in a TAMP problem $\langle s_{0},{\cal A},\mathcal{I},L\rangle$, where $L$ is a text goal description and $\mathcal{I}$ is a set of initial image observations of the scene (see Section 3). It first grounds the set of actions $A$ reachable from $s_{0}$ using ground-actions. Then, it accumulates the set of reachable literals $\Lambda$ by taking the effects of all actions $A$. These sets can be filtered by action or predicate type if it is desired to focus VLM assistance on specific aspects of the planning problem. Then, it prompts query-vlm for a partial plan $a_{1},...,a_{n},l_{1},...,l_{k}]$ using actions $a_{i}\in A$ and goal literals $l_{j}\in\Lambda_{m}$ that achieve the goal description $g$. Importantly, we have the VLM fill in the description parameter $d$ for each of these actions. We then transform the original TAMP problem to force solutions to admit the partial plan as a subsequence. Specifically, we create a predicate Executed that models whether the $i$th action in the plan was executed and add Executed to the effects of action $a_{i}$ and the preconditions of action $a_{i+1}$. Finally, we make the planning goal be $G=\{l_{i},...,l_{m}\}\subseteq L$ and $\textsc{Executed}(n)$, which indicates that all actions have been executed and solve the transformed TAMP problem with a generic TAMP algorithm.
A.1.3 OWL-TAMP Constraint Generation Example Walkthrough
We walk through a concrete example of generating both discrete and continuous constraints for the ‘Mug1’ task (Figure 3a), where the natural language goal is $L$ = “Setup the mug so it’s upright, then put whatever object that fits inside of it”.
First, we prompt the VLM to generate discrete constraints given the task goal. It generates the following partial plan, along with language descriptions filled in for each action:
1(’Pick(mug)’, ’grasp the mug securely to lift it from the table.’)
3(’Pick(fork)’, ’grasp the fork securely to lift it from the table.’)
Next, we prompt the VLM to provide continuous constraints for the achieve_goal operator at the end of the plan. This corresponds to providing continuous constraints that correspond to the task’s goal description.
1# Goal Check Functions:
2def goal_check0() -> bool:
4 return upright_mug
5
6def goal_check1() -> bool:
8 return position_within_bounds(fork.pose, inside_mug_bounds)
We then prompt the VLM to generate continuous constraints for each operator given these generated constraint functions as input. We start from the beginning of the plan. The Pick(mug) action (which corresponds to attach) does not have any effects or constraint predicates that depend on the VLM, so we skip this action. The Place_Ontop(mug, table) action ((which corresponds to detach) does feature a VLMPose constraint. So we prompt the VLM to generate continuous constraints that implement it based on the language description as well as the goal generated constraints.
1# Goal Check Functions:
2def goal_check2() -> bool:
The Place_Inside(fork, mug) action (which also corresponds to detach) does not have any effects or constraint predicates that depend on the VLM, so we skip this action as well. Finally, the Place_Inside(fork, mug) action does have a VLMPose constraint, so we once again prompt the VLM, and obtain the following continuous constraints for this action:
1# Goal Check Functions:
2def goal_check2() -> bool:
4 return position_within_bounds(fork.pose, inside_mug_bounds)
Note that in this particular case, the VLM only checks that the fork is inside the mug, which is a constraint that is built into the TAMP system itself (via the Inside predicate that is an effect of the Place_Inside action).
Now, we have a partial plan along with continuous constraints for particular actions. We call our underlying TAMP system to attempt to satisfy these, as well as its own internal constraints (Appendix A.1.1).
A.2 Additional RAVENS-YCB Manipulation Experiment Details
We provide additional details on the tasks we ran experiments on in the simulated RAVENS-YCB Manipulation environment (Section 6) and discuss the Code as Policies baseline implementation in Appendix A.2.1.
A visualization of an initial state in each of our tasks is shown in Figure 5. In each of these tasks, the robot had access to three parameterized actions: Pick($o,g$) where $o$ is an object and $g$ is a 6-DOF grasp pose consisting of $x,y,z,{\it roll},{\it pitch},{\it yaw}]$ in the world coordinate frame, Place(o,s,p)(o,s,p) where $o$ is the object to be placed, $s$ is the surface or container to place atop or inside, and $p$ is the 6D placement pose at which the robot will move and simply open its gripper, and Pour(o,s,t)(o,s,t), where $o$ is the object to be poured from, $s$ is the surface or container to pour onto or into, and $t$ is a 4D vector consisting of a 3D position and a pitch angle at which to tip the hand to accomplish the pour. For each random seed, the positions of all objects except the ‘table_surface’ object in the scene are randomized unless otherwise stated.
  • $\bullet$
    Berry1: Goal description: “put the strawberry onto the light-grey region at the center of the table”. Objects: strawberry, light_grey_region, table_surface. The position of the light_grey_region is held constant at the center of the table. The optimal solution to this task requires a sequence of 2 actions.
  • $\bullet$
    Citrus: Goal description: “pack the citrus fruit onto the plate”. Objects: strawberry, apple, pear, lemon, orange, plum, table_surface. The plate object is always set to be at the center of the table in the initial state. The optimal solution to this task requires a sequence of 4 actions.
  • $\bullet$
    Berry2: Goal description: “put the strawberry onto the light-grey region at the center of the table”. Objects: strawberry, light_grey_region, potted_meat_can, table_surface. The only object whose position is randomized in the initial state is the strawberry: the potted_meat_can is always set to totally obstruct the surface of the light_grey_region in the center of the table: it must be moved out of the way to make room to place the strawberry and successfully accomplish the goal description. The optimal solution to this task requires a sequence of 4 actions.
  • $\bullet$
    BerryCook: Goal description: “Cook the strawberry by putting it in the pan, then finally simply place it in the bowl. The strawberry should only be in the bowl at the end!”. Objects: strawberry, skillet, bowl, table_surface. The only object whose position is randomized in the initial state is the strawberry. The optimal solution to this task requires a sequence of 4 actions.
  • $\bullet$
    FruitSort: Goal description: “Put all the fruit to the left of the line bisecting the table”. Objects: pear, sugar_box, hammer, tomato_soup_can, strawberry, apple, red_line, table_surface. The position of the red_line object is kept constant at the center of the table. The optimal solution to this task requires a sequence of 6 actions.
  • $\bullet$
    Coffee: Goal description: “I want to pour some coffee into the cup; can you set up the cup on the table so I can do this properly?” (requires reorienting the cup so that it is placed ‘right-side-up’ on the table). Objects: mug, table_surface. The optimal solution to this task requires a sequence of 2 actions.
  • $\bullet$
    Mug1: Goal description: “Setup the mug so it’s upright, then put whatever object that fits inside of it”. Objects: fork, power_drill, potted_meat_can, mug, table_surface. The initial position of the mug is fixed. The optimal solution to this task requires a sequence of 4 actions.
  • $\bullet$
    Mug2: Goal description: “Place cutlery inside the mug and then place the mug itself on the table near the condiment” (the mug’s opening is obstructed by a large orange, which must be moved out of the way). Objects: fork, knife, sponge, strawberry, orange, mustard_bottle, mug, table_surface. The initial position of the mug and orange are fixed such that the orange is always obstructing the mouth of the mug. The optimal solution to this task requires a sequence of 8 actions.
  • $\bullet$
    Mug3: Goal description: “place cutlery into the mug and have the mug be near the condiment”. This is the same as CutleryInMug2, except in the initial state the mug is not obstructed by an orange, but rather has a ball stuck inside it. This must be removed by ‘pouring’ it out of the cup itself). Objects: fork, golf_ball, mustard_bottle, mug, table_surface. The initial position of the mug and golf_ball are fixed such that the golf_ball is always inside the mug. The optimal solution to this task requires a sequence of 8 actions.
  • $\bullet$
    SoupPour: Goal description: “Serve the fruits on the white mat (make sure the peach is to the right of the apple” and pour soup into the red container”. The white mat is originally obstructed by the soup can as well as a spam can, which must be moved out of the way to successfully place the fruits. Objects: white_mat, tomato_soup_can, potted_meat_can, bowl, apple, peach, table_surface. The only object positions that are randomized in the initial state are that of the apple and the peach. The optimal solution to this task requires a sequence of 10 actions.
A.2.1 Code as Policies Implementation Details
Following recent work [10], we adapt Code as Policies to our RAVENS YCB-Manipulation domain by having it output a Language Model Program (LMP) function that — given an initial state — directly outputs a full plan with discrete actions and continuous parameters. We provide it with all the same helper functions used for continuous parameter generation that our method has access to. We also provide it with few-shot example solutions to the same 3 example problems as we provide in the continuous constraint generation prompt of our approach. Additionally, we provide it with one sampler per controller that it can use to generate continuous parameters for each respective controller. We provide the full detailed prompt we use for this method below in Appendix A.8.
Tasks
Table 2: Number of samples on all tasks. We report the average number of continuous parameter sampling attempts for each task across 1010 random seeds; the number of samples within the 95% confidence interval is reported after the $\pm$ sign for every entry. We also include a row (‘Manual’) reporting results on a variant of OWL-TAMP where we manually specify ground-truth discrete and continuous constraints to serve as an approximate lower bound on the number of samples our approach could achieve.
A.3 Additional Real-World Experiment Details
We attempted the following goal descriptions on the real robot:
@@ A.3 Additional Real-World Experiment Details
Please see our supplementary video to see the real robot execution.
A.4 Additional Experimental Results